OpenStudy (anonymous):

Find f^-1(x) if f(x) = -4x - 2.

OpenStudy (anonymous):

According to my tutor, the final answer is 4x - 2

OpenStudy (anonymous):

I believe she is wrong, could someone verify this?

OpenStudy (anonymous):

@Zarkon @kropot72 @goformit100 @ghazi

hartnn (hartnn):

yeah, its incorrect.

hartnn (hartnn):

let y= -4x-2 to get inverse function, we first replace 'x' by 'y' and 'y' by 'x' so, x = -4y-2 got this ? now try to isolate (find) y from here.

OpenStudy (anonymous):

x=-2?

OpenStudy (anonymous):

So would it be x-2/4 ?

hartnn (hartnn):

x = just -2 ? x = -4y -2 add 2 on both sides, what u get ?

OpenStudy (anonymous):

I got 2x=-4y

hartnn (hartnn):

no no ... x = -4y -2 x+2 = -4y -2+2 so, x+2 = -4y got this ?

OpenStudy (anonymous):

Got it

hartnn (hartnn):

now, x+2 = -4y so, multiplying -1 on both sides -x-2 = 4y just divide both sides by 4 and you'll get y

OpenStudy (anonymous):

y= -1/2

hartnn (hartnn):

you just won't get value of y you'll get a function of 'x' (-x-2)/4 = y thats your f^-1(x) = -(x+2)/4

OpenStudy (anonymous):

You're the best at explaining things @hartnn thankss!

hartnn (hartnn):

hey thanks! :) and welcome ^_^